OpenStudy (anonymous):

If cos(theta)=-4/5 and theta lies in Quadrant 2, what is the value of tan (theta)

OpenStudy (ash2326):

Could you tell me if cos theta= 4/5, then what's tan theta ? We'll ignore the "-" sign, then we'll come back to it

OpenStudy (anonymous):

umm 5/4

OpenStudy (ash2326):

How did you get this?

OpenStudy (anonymous):

i drew the triangle on the unit circle and marked cos theta and labeled the sides with the numbers and than found tan theta

OpenStudy (ash2326):

 The third side will be \[\sqrt{5^2-4^2}=3\]

OpenStudy (ash2326):